# Physics-inspired Generative Models in Medical Imaging: A Review

### v1

@@ Abstract
Physics-inspired generative models, in particular diffusion and Poisson flow models, enhance Bayesian methods and promise great utilities in medical imaging. This review examines the transformative role of such generative methods. First, a variety of physics-inspired generative models, including Denoising Diffusion Probabilistic Models (DDPM), Score-based Diffusion Models, and Poisson Flow Generative Models (PFGM and PFGM++), are revisited, with an emphasis on their accuracy, robustness as well as acceleration. Then, major applications of physics-inspired generative models in medical imaging are presented, comprising image reconstruction, image generation, and image analysis. Finally, future research directions are brainstormed, including unification of physics-inspired generative models, integration with vision-language models (VLMs),and potential novel applications of generative models. Since the development of generative methods has been rapid, this review will hopefully give peers and learners a timely snapshot of this new family of physics-driven generative models and help capitalize their enormous potential for medical imaging.
@@ 1 INTRODUCTION
In retrospect, variational autoencoders (VAEs) and generative adversarial networks (GANs) were among the first to pave the way. VAEs work by encoding input data into a lower-dimensional space and then synthesize samples in the original space, obeying the distribution of the original data. GANs, on the other hand, consist of two competing networks: a generator that creates data and a discriminator that separates generated and real data. The two networks interact to define the adversarial process that improves the quality of the generated data so that they are indistinguishable from the original data. More recently, physics-inspired generative models have become the state of the art, with the diffusion models and Poisson flow generative models being the most popular, which construct an image or sample by starting from a random distribution and refining it towards the distribution of the target data through a sampling process in a number of steps (from thousands in the original setting to one in the consistency model). These physics-inspired generative models can generate realistic, diverse, and complex data in a principled yet universal way, already demonstrating a huge potential for medical imaging. Notably, as we will see in this review, VAEs are making a silent comeback as a key constructplaying a pivotal role in the state of the art generative methods.
@@
An overview of diffusion models was first posted in November 2022, revised in June 2023, and formally published in Aug 2023 in Medical Image Analysis [54]. It cited 192 key papers and provides an in-depth coverage of the theoretical foundations of diffusion models, their practical applications in medical imaging using the taxonomy built on nine sub-fields such as image translation, reconstruction, and anomaly detection, as well as the challenges and future directions. The survey categorized diffusion models based on their applications, imaging modality, organ of interest, and algorithms, offering a multi-perspective categorization and highlighting the importance of these models in generating synthetic medical images. Such realistically simulated samples can alleviate data scarcity and improve network training for superior performance.
@@
Another interesting review published in August 2023 analyzed 80 papers on image generation using diffusion models [118]. The difficulties were examined in generating images with multiple objects, creating images of rare/novel objects and improving quality of generated images. The paper organized the discussion around these challenges and presented various methods to address them, such as leveraging layout information, refining attention maps, and employing retrieval-based methods. It also touched upon the importance of text encoders in improving image quality and explored the use of mixtures of experts and reinforcement learning to enhance outcomes.
@@
Another survey over 160 papers was posted in 2022 and published in IEEE Transactions on Pattern Analysis and Machine Intelligence (TPAMI) in September 2023 [27]. It covered diffusion models in the field of computer vision and addressed the three schemes: Denoising Diffusion Probabilistic Models (DDPM), Noise Conditioned Score Networks (NCSN), and score-based diffusion models via stochastic differential equations (SDEs). The paper also discussed the relationship between diffusion models and other generative models, presenting a categorization of diffusion models applied in computer vision and identifying research directions, particularly the need to improve sampling/inference efficiency without compromising the quality of generated samples. In yet another comprehensive review of 258 papers on diffusion models in various domains, the core aspects were presented of the formulation, algorithms, and applications [12]. Through gradual refinement of the review from September 2022 (version 1) to December 2023 (version 10), this review highlighted methods for generating high-fidelity samples across different modalities and discussed strategies for sampling acceleration, diffusion process design, and likelihood optimization. The article was concluded with discussions on the current limitations and future advancements.
@@
While each review had different methodological emphases and covered various spectra of applications, they have collectively underscored the significance of diffusion models across different fields, with the common core challenge to optimize the image quality and sampling efficiency individually or simultaneously.
@@
While the above reviews/surveys are excellent and complementary, the rationale for us to write this new review is still compelling. Recently, especially over the past year, the emergence of Poisson flow generative models [108, 109] represents a significant leap forward in the field of generative AI, modeling complex data distributions and producing high-fidelity samples with the state-of-the-art performance. Like diffusion models, Poisson flow or other physics-/mathematics-inspired generative models can be cast in the score-matching format. Unlike diffusion models, which draw their inspiration from non-equilibrium thermodynamics, Poisson flow generative models are grounded in the principle of electrostatics, bringing new theoretical insights and empowering technical capabilities. This foundational shift offers a distinct advantage in terms of precision and efficiency. Poisson equations allow for a more direct manipulation of the data distribution, enabling a better controlled and potentially faster convergence to the target distribution. At the same time, impressive new results have been reported on the classic diffusion models such as consistency models. Therefore, the benefits to thoroughly summarize and analyze these advancements through a new review article are clear; mainly, to clarify the new thinking and mechanisms, capitalize the theoretical and technical potentials, and facilitate real-world applications of contemporary generative AI.
@@
The rest of this review is organized as follows. In the second section, physics-driven generative models are critically described from diffusion models to Poisson flow models, along with cutting-edge techniques for accelerated inference. In the third section, medical imaging applications of these generative models are concisely summarized, involving image reconstruction, image generation, and image analysis. In the last section, future directions are openly discussed, as related to unification of physics-inspired generative models, integration with vision-language models (VLMs),and potential novel applications of generative models.
@@ 2 PHYSICS-INSPIRED GENERATIVE MODELS
In this section, we give an overview of physics-inspired generative models, going from the earlier versions of diffusion models based on Markov chains all the way to the latest developments employing a combination of a score matching loss, distillation loss, and an adversarial loss to achieve state-of-the-art image quality and sampling efficiency. In doing so, we will place an particular emphasis on the generalized framework referred to as Poisson Flow Generative Models (PFGM)++ [109], as this takes a key step towards the unification of physics-inspired generative models.
2.1 Denoising Diffusion Probabilistic Models (DDPM)
Denosing Diffusion Probabilistic Models (DDPM)  [78, 46] are a class of generative models that have gained significant attention for their ability to generate high-quality images without adversarial training. The core idea behind DDPM is to model the generation process as an iterative denoising process over a sequence of time steps, starting from pure noise and progressively adding structure to reach a sample from the data distribution. More formally, for each training data point $\bm{x}_{0}\sim p(\bm{x})$ a discrete Markov chain ${\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{N}}$ is constructed with transition kernel $p(\bm{x}_{i}|\bm{x}_{i-1})=\mathcal{N}(\bm{x}_{i};\sqrt{\alpha_{i}}\bm{x}_{i},(1-\alpha_{i})\bm{I})$ for $\alpha_{i}:=\prod_{j=1}^{i}(1-\beta_{j}).$ A sequence of positive noise scales $10<\beta_{1},\beta_{2},...,\beta_{N}<$ 1 will be chosen such that $\bm{x}_{N}$ are approximately distributed as $\mathcal{N}(\bm{0},\bm{I})$. The reverse variational Markov chain is parameterized as $p_{\bm{\theta}}(\bm{x}_{i-1}|\bm{x}_{i})=\mathcal{N}(\bm{x}_{i-1};\frac{1}{\sqrt{1-\beta_{i}}}(\bm{x}_{i}+\beta_{i}s_{\bm{\theta}}(\bm{x}_{i},i),\beta_{i}\bm{I}))$ and is learned through the objective:
@@ 2.1 Denoising Diffusion Probabilistic Models (DDPM)
starting from $\bm{x}_{N}\sim\mathcal{N}(\bm{0},\bm{I})$. The key to the success of DDPM is that they transform the problem of generating samples from complex data distributions into a sequence of simpler denoising tasks, enabling high-quality image generation without adversarial training.
2.2 Noise Conditional Score Networks (NCSN)
@@
2.3 Score-based Diffusion Models
Score-based diffusion models [84], also known as score-based generative models, provides a unified framework for score-based models [82] and diffusion models [78, 46] by formulating the diffusion process as a stochastic differential equation (SDE). Intuitively, this takes the idea of perturbing the data and learning to model complex data distributions via a sequence of simpler denoising task to the limit where the number of noise scales becomes infinite, leveraging advances in score modelling to obtain an effective sequential denoiser. This SDE takes can be written on the general form
@@ 2.3 Score-based Diffusion Models
As with NSCN, the score function estimated via a DSM objective. Notably, as stated above, this frameworks includes DDPM and NCSN as special cases as they can be regarded as discretizations of two particular SDEs.
@@
Once equipped with this estimate, we can generate an image by drawing an initial sample from the prior noise distribution and solving Equation 8 using some numeric ODE solver. Score-based diffusion models offer a powerful and elegant framework for generative modeling, leveraging the mathematics of diffusion processes and score functions to transform noise into high fidelity samples.
Figure 2: Physics-inspired deep generative models. (a) Schematic of forward and reverse SDE on score based diffusion models [84]; (b) Overview of PFGM++ [109]; (c) Schematic overview of LDM [71]; (d) Schematic of mapping learned by consistency models [85]. Note that different notation may have been used in the text compared to this figure.
2.4 Poisson Flow Generative Models (PFGM)
Poisson Flow Generative Models (PFGM) [108] are a relatively new family of deep generative models inspired by electrostatics, instead of non-equilibrium thermodynamics as is the case for diffusion models. In particular, PFGM treats the $N$-dimensional data $\bm{x}\in\mathbb{R}^{N}$, as electric charges in a $N+1$-dimensional augmented space $\bm{\tilde{x}}:=(\bm{x},z)\in\mathbb{R}^{N+1}$. Following the law of motion defined by the electric field line emitted by the charges, the data distribution placed on the $z=0$ hyperplane can be transformed into a distribution on the hemisphere with radius $r.$ As $r\rightarrow\infty$ this distribution will become uniform [108]. Instead of estimating a time-dependent score function, as in score-based diffusion models, PFGM learns an empirical Poisson field, derived from the Poisson equation in the augmented space,
@@ 2.4 Poisson Flow Generative Models (PFGM)
where $p(\bm{y})$ is the data distribution and $S_{N}(1)$ denotes the surface area of a unit N-sphere. We can move from the data distribution to the uniform distribution on the infinite hemisphere by running the ODE $d\bm{\tilde{x}}=\bm{E}(\bm{\tilde{x}})dt$ forward and back in time. In practice, PFGM is trained using a large batch of size $n$ to approximate the integral in Equation 10 to estimate a normalized empirical Poisson field $\bm{\hat{E}}(\bm{\tilde{x}})=c(\bm{\tilde{x}})\sum_{i=1}^{n}\frac{\bm{\tilde{x}}-\bm{\tilde{y}}_{i}}{||\bm{\tilde{x}}-\bm{\tilde{y}}_{i}||^{N+1}}.$ The training objective is to minimize the $\ell_{2}$-loss between the network $f_{\theta}(\bm{\tilde{x}})$ and $\bm{E}(\bm{\tilde{x}})/||\bm{E}(\bm{\tilde{x}})||$ over various, carefully selected to ensure coverage of sampling trajectories, positions of $\bm{\tilde{x}}.$ PFGM represent a powerful approach to generative modeling, combining the stochastic modeling capabilities of Poisson processes with flow-based generative techniques.
@@ 2.5 PFGM++
PFGM++ [109] takes an important step towards unification of physics-inspired generative models since it contains PFGM [108] and, remarkably, score-based diffusion models [84, 52] as special cases. An overview is available in Figure 2 (b). PFGM++ in essence generalize PFGM in terms of augmented space dimensionality. The $N$-dimensional data, $\bm{x}\in\mathbb{R}^{N}$, are treated as positive electric charges in a $N+D$-dimensional augmented space. Let $\tilde{\bm{x}}:=(\bm{x},\bm{z})\in\mathbb{R}^{N+D}$ and $\tilde{\bm{y}}:=(\bm{x},\bm{0})\in\mathbb{R}^{N+D}$ denote the augmented perturbed and ground truth data, respectively. The electric field lines generated by these charges define a surjection via the ODE, $d\bm{\tilde{x}}=\bm{E}(\bm{\tilde{x}})dt$, between an uniform distribution on the infinite $N+D$-dimensional hemisphere and the data placed on the $\bm{z}=\bm{0}$ hyperplane. The objective of interest in PFGM++ is the high dimensional electric field
@@
where $S_{N+D-1}(1)$ is the surface area of the unit $(N+D-1)$-sphere, and $p(\bm{y})$ the ground truth data distribution. Xu et al. [109] notes that the electric field is rotationally symmetric on the $D$-dimensional cylinder $\sum_{i=1}^{D}z_{i}^{2}=r^{2},\forall r>0$ and it suffices to track the norm of the augmented variables $r(\bm{\tilde{x}}):=||\bm{z}||_{2}.$ In particular, we can observe that
@@
2.5.1 PFGM++ and score-based diffusion models in theory and in practice
PFGM++ admits PFGM [108] ($D=1$) and, remarkably, score-based diffusion models [84, 52] ($D\to\infty$) as special cases. In addition to being intimately connected in theory, PFGM++ and score-based diffusion models [84, 52] are also closely connected in practice. Xu et al. [109] demonstrates that one can re-purpose the training and sampling algorithms developed for diffusion models in [52] to PFGM++. This follows from the fact that $\sigma:=\sigma(t)=t$ in [52], $\bm{\tilde{x}}:=(\bm{x},r)$, and the alignment formula $r=\sigma\sqrt{D}$, which allows for the change-of-variables $d\bm{x}=f^{*}_{\theta}(\bm{\tilde{x}})/\sqrt{D}dr=f^{*}_{\theta}(\bm{\tilde{x}})dt$ since $dr=d\sigma\sqrt{D}=dt\sqrt{D}.$

### v2

@@ Abstract
Physics-inspired Generative Models (GMs), in particular Diffusion Models (DMs) and Poisson Flow Models (PFMs), enhance Bayesian methods and promise great utility in medical imaging. This review examines the transformative role of such generative methods. First, a variety of physics-inspired GMs, including Denoising Diffusion Probabilistic Models (DDPMs), Score-based Diffusion Models (SDMs), and Poisson Flow Generative Models (PFGMs and PFGM++), are revisited, with an emphasis on their accuracy, robustness as well as acceleration. Then, major applications of physics-inspired GMs in medical imaging are presented, comprising image reconstruction, image generation, and image analysis. Finally, future research directions are brainstormed, including unification of physics-inspired GMs, integration with Vision-Language Models (VLMs), and potential novel applications of GMs. Since the development of generative methods has been rapid, this review will hopefully give peers and learners a timely snapshot of this new family of physics-driven generative models and help capitalize their enormous potential for medical imaging.
@@ 1 INTRODUCTION
In retrospect, Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) were among the first to pave the way. VAEs work by encoding input data into a lower-dimensional space and then synthesize samples in the original space, obeying the distribution of the original data. GANs, on the other hand, consist of two competing networks: a generator that creates data and a discriminator that separates generated and real data. The two networks interact to define the adversarial process that improves the quality of the generated data so that they are indistinguishable from the original data. More recently, physics-inspired Generative Models (GMs) have become the state of the art, with the Diffusion Models (DMs) and Poisson Flow Generative Models (PGFMs) being the most popular ones. They construct an image or sample by starting from a random distribution and refining it towards the distribution of the target data through a sampling process in a number of steps (from thousands in the original setting to one in the consistency model). These physics-inspired GMs can generate realistic, diverse, and complex data in a principled yet universal way, already demonstrating a huge potential for medical imaging. Notably, as we will see in this review, VAEs are making a silent comeback as a key construct playing a pivotal role in the state-of-the-art generative methods.
@@
An overview of DMs was first posted in November 2022, revised in June 2023, and formally published in Aug 2023 in Medical Image Analysis [56]. It cited 192 key papers and provides an in-depth coverage of the theoretical foundations of DMs, their practical applications in medical imaging using the taxonomy built on nine sub-fields such as image translation, reconstruction, and anomaly detection, as well as the challenges and future directions. The survey categorized DMs based on their applications, imaging modality, organ of interest, and algorithms, offering a multi-perspective categorization and highlighting the importance of these models in generating synthetic medical images. Such realistically simulated samples can alleviate data scarcity and improve network training for superior performance.
@@
Another interesting review published in August 2023 analyzed 80 papers on image generation using DMs [124]. The difficulties were examined in generating images with multiple objects, creating images of rare/novel objects and improving quality of generated images. The paper organized the discussion around these challenges and presented various methods to address them, such as leveraging layout information, refining attention maps, and employing retrieval-based methods. It also touched upon the importance of text encoders in improving image quality and explored the use of mixtures of experts and reinforcement learning to enhance outcomes.
@@
Another survey over 160 papers was posted in 2022 and published in IEEE Transactions on Pattern Analysis and Machine Intelligence (TPAMI) in September 2023 [28]. It covered DMs in the field of computer vision and addressed the three schemes: Denoising Diffusion Probabilistic Models (DDPMs), Noise Conditioned Score Networks (NCSNs), and Score-based Diffusion Models (SDMs) via stochastic differential equations (SDEs). The paper also discussed the relationship between DMs and other GMs, presenting a categorization of DMs applied in computer vision and identifying research directions, particularly the need to improve sampling/inference efficiency without compromising the quality of generated samples. In yet another comprehensive review of 258 papers on DMs in various domains, the core aspects were presented of the formulation, algorithms, and applications [13]. Through gradual refinement of the review from September 2022 (version 1) to December 2023 (version 10), this review highlighted methods for generating high-fidelity samples across different modalities and discussed strategies for sampling acceleration, diffusion process design, and likelihood optimization. The article was concluded with discussions on the current limitations and future advancements.
@@
While each review had different methodological emphases and covered various spectra of applications, they have collectively underscored the significance of DMs across different fields, with the common core challenge to optimize the image quality and sampling efficiency individually or simultaneously.
@@
While the above reviews/surveys are excellent and complementary, the rationale for us to write this new review is still compelling. Recently, especially over the past year, the emergence of PFGMs [113, 114] represents a significant leap forward in the field of generative AI, modeling complex data distributions and producing high-fidelity samples with the state-of-the-art performance. Like DMs, Poisson flow or other physics-/mathematics-inspired GMs can be cast in the score-matching format. Unlike DMs, which draw their inspiration from non-equilibrium thermodynamics, PFGM are grounded in the principle of electrostatics, bringing new theoretical insights and empowering technical capabilities. This foundational shift offers a distinct advantage in terms of precision and efficiency. Poisson equations allow for a more direct manipulation of the data distribution, enabling a better controlled and potentially faster convergence to the target distribution. At the same time, impressive new results have been reported on the classic DMs such as consistency models. Therefore, the benefits to thoroughly summarize and analyze these advancements through a new review article are clear; mainly, to clarify the new thinking and mechanisms, capitalize the theoretical and technical potentials, and facilitate real-world applications of contemporary generative AI.
@@
The rest of this review is organized as follows. In the second section, physics-driven GMs are critically described from DMs to PFGM++, along with cutting-edge techniques for accelerated inference. In the third section, medical imaging applications of these GMs are concisely summarized, involving image reconstruction, image generation, and image analysis. In the last section, future directions are openly discussed, as related to unification of physics-inspired GMs, integration with VLMs, and potential novel applications of GMs.
@@ 2 PHYSICS-INSPIRED GENERATIVE MODELS
In this section, we give an overview of physics-inspired GMs, going from the earlier versions of DMs based on Markov chains all the way to the latest developments employing a combination of a score matching loss, distillation loss, and an adversarial loss to achieve state-of-the-art image quality and sampling efficiency. In doing so, we will place an particular emphasis on the generalized framework referred to as PFGM++ [114], as this takes a key step towards the unification of physics-inspired GMs.
2.1 DDPMs
DDPMs  [83, 48] are a class of GMs that have gained significant attention for their ability to generate high-quality images without adversarial training. The core idea behind DDPMs is to model the generation process as an iterative denoising process over a sequence of time steps, starting from pure noise and progressively adding structure to reach a sample from the data distribution. More formally, for each training data point $\bm{x}_{0}\sim p(\bm{x})$ a discrete Markov chain ${\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{N}}$ is constructed with transition kernel $p(\bm{x}_{i}|\bm{x}_{i-1})=\mathcal{N}(\bm{x}_{i};\sqrt{\alpha_{i}}\bm{x}_{i},(1-\alpha_{i})\bm{I})$ for $\alpha_{i}:=\prod_{j=1}^{i}(1-\beta_{j}).$ A sequence of positive noise scales $10<\beta_{1},\beta_{2},...,\beta_{N}<$ 1 will be chosen such that $\bm{x}_{N}$ are approximately distributed as $\mathcal{N}(\bm{0},\bm{I})$. The reverse variational Markov chain is parameterized as $p_{\bm{\theta}}(\bm{x}_{i-1}|\bm{x}_{i})=\mathcal{N}(\bm{x}_{i-1};\frac{1}{\sqrt{1-\beta_{i}}}(\bm{x}_{i}+\beta_{i}s_{\bm{\theta}}(\bm{x}_{i},i),\beta_{i}\bm{I}))$ and is learned through the objective:
@@ 2.1 DDPMs
starting from $\bm{x}_{N}\sim\mathcal{N}(\bm{0},\bm{I})$. The key to the success of DDPMs is that they transform the problem of generating samples from complex data distributions into a sequence of simpler denoising tasks, enabling high-quality image generation without adversarial training.
2.2 NCSNs
@@
2.3 SDMs
SDMs [89], also known as score-based GMs, provide a unified framework for score-based models [87] and DMs [83, 48] by formulating the diffusion process as a stochastic differential equation (SDE). Intuitively, this takes the idea of perturbing the data and learning to model complex data distributions via a sequence of simpler denoising task to the limit where the number of noise scales becomes infinite, leveraging advances in score modelling to obtain an effective sequential denoiser. This SDE takes can be written on the general form
@@ 2.3 SDMs
As with NSCNs, the score function estimated via a DSM objective. Notably, as stated above, this frameworks includes DDPM and NCSN as special cases as they can be regarded as discretizations of two particular SDEs.
@@
Once equipped with this estimate, we can generate an image by drawing an initial sample from the prior noise distribution and solving Equation 8 using some numeric ODE solver. SDMs offer a powerful and elegant framework for generative modeling, leveraging the mathematics of diffusion processes and score functions to transform noise into high fidelity samples.
Figure 2: Physics-inspired deep GMs. (a) Schematic of forward and reverse SDE on SDMs [89]; (b) Overview of PFGM++ [114]; (c) Schematic overview of LDM [76]; (d) Schematic of mapping learned by consistency models [90]. Note that different notation may have been used in the text compared to this figure.
2.4 PFGMs
PFGMs [113] are a relatively new family of deep GMs inspired by electrostatics, instead of non-equilibrium thermodynamics as is the case for DMs. In particular, PFGMs treat the $N$-dimensional data $\bm{x}\in\mathbb{R}^{N}$, as electric charges in a $N+1$-dimensional augmented space $\bm{\tilde{x}}:=(\bm{x},z)\in\mathbb{R}^{N+1}$. Following the law of motion defined by the electric field line emitted by the charges, the data distribution placed on the $z=0$ hyperplane can be transformed into a distribution on the hemisphere with radius $r.$ As $r\rightarrow\infty$ this distribution will become uniform [113]. Instead of estimating a time-dependent score function, as in SDMs, PFGMs learn an empirical Poisson field, derived from the Poisson equation in the augmented space,
@@ 2.4 PFGMs
where $p(\bm{y})$ is the data distribution and $S_{N}(1)$ denotes the surface area of a unit N-sphere. We can move from the data distribution to the uniform distribution on the infinite hemisphere by running the ODE $d\bm{\tilde{x}}=\bm{E}(\bm{\tilde{x}})dt$ forward and back in time. In practice, PFGMs are trained using a large batch of size $n$ to approximate the integral in Equation 10 to estimate a normalized empirical Poisson field $\bm{\hat{E}}(\bm{\tilde{x}})=c(\bm{\tilde{x}})\sum_{i=1}^{n}\frac{\bm{\tilde{x}}-\bm{\tilde{y}}_{i}}{||\bm{\tilde{x}}-\bm{\tilde{y}}_{i}||^{N+1}}.$ The training objective is to minimize the $\ell_{2}$-loss between the network $f_{\theta}(\bm{\tilde{x}})$ and $\bm{E}(\bm{\tilde{x}})/||\bm{E}(\bm{\tilde{x}})||$ over various, carefully selected to ensure coverage of sampling trajectories, positions of $\bm{\tilde{x}}.$ PFGMs represent a powerful approach to generative modeling, combining the stochastic modeling capabilities of Poisson processes with flow-based generative techniques.
@@ 2.5 PFGM++
PFGM++ [114] take an important step towards unification of physics-inspired GMs since it contains PFGMs [113] and, remarkably, SDMs [89, 54] as special cases. An overview is available in Figure 2 (b). PFGM++ in essence generalize PFGMs in terms of augmented space dimensionality. The $N$-dimensional data, $\bm{x}\in\mathbb{R}^{N}$, are treated as positive electric charges in a $N+D$-dimensional augmented space. Let $\tilde{\bm{x}}:=(\bm{x},\bm{z})\in\mathbb{R}^{N+D}$ and $\tilde{\bm{y}}:=(\bm{x},\bm{0})\in\mathbb{R}^{N+D}$ denote the augmented perturbed and ground truth data, respectively. The electric field lines generated by these charges define a surjection via the ODE, $d\bm{\tilde{x}}=\bm{E}(\bm{\tilde{x}})dt$, between an uniform distribution on the infinite $N+D$-dimensional hemisphere and the data placed on the $\bm{z}=\bm{0}$ hyperplane. The objective of interest in PFGM++ is the high dimensional electric field
@@
where $S_{N+D-1}(1)$ is the surface area of the unit $(N+D-1)$-sphere, and $p(\bm{y})$ the ground truth data distribution. Xu et al. [114] note that the electric field is rotationally symmetric on the $D$-dimensional cylinder $\sum_{i=1}^{D}z_{i}^{2}=r^{2},\forall r>0$ and it suffices to track the norm of the augmented variables $r(\bm{\tilde{x}}):=||\bm{z}||_{2}.$ In particular, we can observe that
@@
2.5.1 PFGM++ and SDMs in theory and in practice
PFGM++ admit PFGMs [113] ($D=1$) and, remarkably, SDMs [89, 54] ($D\to\infty$) as special cases. In addition to being intimately connected in theory, PFGM++ and SDMs [89, 54] are also closely connected in practice. Xu et al. [114] demonstrate that one can re-purpose the training and sampling algorithms developed for DMs in [54] to PFGM++. This follows from the fact that $\sigma:=\sigma(t)=t$ in [54], $\bm{\tilde{x}}:=(\bm{x},r)$, and the alignment formula $r=\sigma\sqrt{D}$, which allows for the change-of-variables $d\bm{x}=f^{*}_{\theta}(\bm{\tilde{x}})/\sqrt{D}dr=f^{*}_{\theta}(\bm{\tilde{x}})dt$ since $dr=d\sigma\sqrt{D}=dt\sqrt{D}.$